# Energy Extraction from a Kerr Black Hole via Magnetic Reconnection within the Plunging Region

### v1

@@ Abstract
Magnetic reconnection within a highly magnetized plasma has been seen as a viable mechanism to extract the energy from a rotating black hole, as it can generate negative energy plasmoids in the ergoregion. For a typical accreting black hole, the ergoregion is filled with bulk plasma plunging from the innermost-stable-circular orbit (ISCO). In this study, we present an analytical study of the energy extraction via magnetic reconnection process in the plunging region. In contrast to the toroidal plasma, where the magnetic field cannot be derived from the MHD scheme, the magnetic field in the plunging plasma was determined by the ideal-MHD condition. We derive the global magnetic field structure in a fast reconnection model, and we read the expressions for the energies of plasmoids ejected from the reconnection region, for general stationary and axisymmetric spacetimes. Then, we demonstrate the behaviors of ejected energies varying with the reconnection locations in the Kerr spacetime, and identify the region where a negative-energy plasmoid can be produced. We find that for a certain magnetization there exists a critical value of the black hole spin, beyond which the energy extraction can occur, and the energy extraction is most efficient for the near-extreme black hole. Moreover, we study the conditions necessary for a plasmoid with positive energy to escape to the infinity, a crucial requirement for effective energy extractions. Considering the escaping conditions, we provide the parameter space in the radius-spin plane in which the energy extraction mechanism is effective.
1School of Physics, Peking University, No.5 Yiheyuan Rd, Beijing 100871, P.R. China
2Center for High Energy Physics, Peking University, No.5 Yiheyuan Rd, Beijing 100871, P. R. China
3 Collaborative Innovation Center of Quantum Matter, No.5 Yiheyuan Rd, Beijing 100871, P. R. China
@@ 1 Introduction
Beyond the above mentioned mechanisms, it has been proposed that the energy extraction is viable via magnetic reconnections. Koide and Arai [8] proposed the generation of negative energy plasmoids in strongly magnetized fluids within the ergosphere through a purely magnetohydrodynamics (MHD) process. Magnetized accretion flows around black holes can liberate electromagnetic energy in the regions where magnetic field lines are highly curved and reconnect to discharge their energy [9, 10, 11, 12]. This energy release propels the plasma into at least two plasmoids, ejected from the reconnection region. If the released magnetic energy is sufficiently strong, a plasmoid can possess sufficiently negative angular momentum and negative energy as viewed from the infinity. Subsequent to this, Comisso and Asenjo [13] calculated the energy extraction efficiency and power, showing that such an energy extraction rate could be comparable to that of the Blandford-Znajek (BZ) process, thereby implying its astrophysical significance. Subsequent studies have explored energy extraction via magnetic reconnections in different types of spacetimes [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24].
@@ 2.1 The bulk plasma
The bulk plasma around the black hole could be approximated as a perfect fluid, whose stress-energy tensor reads
@@
where $e=-\alpha T^{t}_{\,\,\,t}$ is called the energy-at-infinity density, and $S^{i}=-\alpha T^{i}_{\,\,\,t}$ is the flux density. In our work, we are interested in a stationary, axisymmetric and geometrically thin fluid configuration. Thus, all the physical quantities are defined over the equatorial plane, neglecting their $\theta$-dependence. The fluid is considered to flow along timelike geodesics such that $U_{\mu}dx^{\mu}=-Edt+U_{r}(r)dr+Ld\phi$, with $E,L$ being the conserved quantities along the streamline.
@@ 2.4 Energies of ejected plasmoids
During the reconnection process, most of the magnetic energy is converted to the plasma kinetic energy in the current sheet. Thus, within an ejected plasmoid, we can neglect the magnetic energy and only consider the fluid part. On one hand, we take the incompressible sphere approximation, under which the proper thermal quantities of plasma keep constant during the reconnection process [8]. Thus, the proper enthalpy density in the plasmoid is equal to that of the bulk plasma at the $X$-point, and the enthalpy reads $\Delta H=\omega\Delta V$, assuming that the plasmoid has a small proper volume $\Delta V$. On the other hand, the fluid energy-at-infinity density is defined in the normal tetrad, taken to be $e=-\alpha T^{t}_{\,\,\,t}=-\alpha(\omega u^{t}u_{t}+p)$, as has been introduced in Eq. (2.5). The volume of the plasmoid measured by the normal tetrad is $\Delta\hat{V}=\Delta V/\hat{u}^{(t)}=\Delta V/(\alpha u^{t})$, and the energy is $\Delta E=e\Delta\hat{V}$. Based on these, we can define the energy-at-infinity per enthalpy for the ejected plasmoids at the $X$-point as
@@
We stress that the above discussion is quite general and applicable to any stationary and axisymmetric spacetime. According to Eq. (2.4), it is clear that the influence of the spacetime is directly encoded in $\alpha$ and $\beta^{\phi}$, while indirectly encoded in the fluid velocity and the magnetic field angle. The value of $v_{out}$ depends on the reconnection model, which has no association with the background spacetime.
@@ 3 Energy extraction from a Kerr geometry
In what follows we only consider the prograde case. [^2] For the retrograde case, we confirm that the ejected plasmoids have $\epsilon_{\pm}>0$, even under strong magnetizations. Thus, energy extractions via magnetic reconnection cannot occur within the retrograde plasma. In the Kerr geometry, the ISCO is the triple root of the effective potential for radial motion, leading to a simple form of the radial velocity of the plunging geodesics [44],
@@ 3.1 The magnetic field angle
Figure 2: Behaviors of the magnetic field angle in the bulk plasma plunging from the ISCO (Solid lines) in Kerr geometry. We also plot the results for the infalling plasma with $E=1,L=0$ as a comparision (Dashed lines) . Left: $\xi_{B}$ as a function of $r$ for different $a$ (0.6, 0.94, 0.998). Right: The value of $\xi_{B}$ at the event horizon $r_{H}=1+\sqrt{1-a^{2}}$, as a function of $a$.
@@ 3.2 The energy extraction efficiency
Figure 4: The local energy extraction efficiency within the plunging plasma. Left: $\eta$ as a function of the reconnection location, for differing values of $a$. The magnetization is fixed at $\sigma_{0}=30$. Right: The maximal value of $\eta$ as a function of the black hole spin, under various magnetizations.
In the left panel of Fig. 4, we depict the behavior of the local energy extraction efficiency as a function of the reconnection location within the ISCO. Our findings indicate that for a strong magnetization scenario where the ISCO is situated inside the ergosphere, i.e., $r_{ms}\leq 2$, the energy extraction is inevitably localized near the ISCO. This observation becomes evident when we consider the limit $\sigma_{0}\gg 1$, which yields $v_{out}\approx 1$, $\gamma_{out}\approx\sqrt{(1-{\sf g}^{2})^{3}\sigma_{0}}/(1+{\sf g}^{2})$ with $\sf g$ $=0.49$ for the high magnetization limit. In the vicinity of the ISCO, we have $U^{r},\hat{v}^{(r)}_{s}\rightarrow 0$. Hence, from Eq. (2.4), Eq. (3.7) we can derive
@@
The right panel of Fig. 4 displays the maximum value of $\eta$ under various magnetizations, which depends on the spin of the black hole. For moderate magnetization, such as $\sigma_{0}=3$, only black holes with a rather high spin permit energy extraction. In the case of extreme magnetization, exemplified by $\sigma_{0}=100$, the local energy extraction efficiency exceeds 1 across the range of $a\gtrsim 0.9$. In all these scenarios, the rate of increase of $\eta_{max}$ with respect to $a$ escalates considerably as $a$ approaches its extreme value. To estimate the maximum value of $\eta_{max}$ for a near-extreme black hole, we focus on the near-ISCO region, where the fluid velocity behaves as
@@
Additionally, we discuss the case involving the infalling bulk plasma in A.1, where we confirm that energy extraction cannot occur within the infalling bulk plasma.
@@ 4 Summary and Discussions
In this study, we conducted an in-depth study of energy extraction from a Kerr black hole via magnetic reconnections occurring in the plunging region. We initiated our study by considering a general stationary and axisymmetric spacetime, assuming that the plasma fluid in the equatorial plane adheres to the Killing symmetry of the spacetime. The accompanied magnetic field was set to reverse its direction near the equatorial plane, a region conducive to potent magnetic reconnection. Utilizing the ideal-MHD condition, we obtained the global structure of the magnetic field, which fixes the ejection directions of the plasmoids. Our following analysis incorporated a fast reconnection model characterized by strong magnetic pressure, wherein electromagnetic energy is swiftly liberated, resulting in the formation of two high-velocity plasmoids moving in opposite directions out of the current sheet. Subsequently, we derived the analytical form of the energy-at-infinity per enthalpy for the ejected plasmoids.
@@
Furthermore, we investigated the conditions necessary for a plasmoid with positive energy to escape to the infinity, a crucial requirement for the completion of the energy extraction. By scrutinizing the equation of motion for the ejected plasmoid, we pinpointed a critical radius of the $X$-point, beyond which the plasmoid with positive energy can successfully escape. By combining the escaping conditions with the requirement for producing a negative energy plasmoid, we determined the region where energy extraction is effective, as shown in the green region in Fig. 6. Broadly speaking, the lower boundary of the effective region is dictated by the critical radius for the escaping, and the upper boundary is determined by the upper limit for negative energy when the ISCO is situated outside the ergosphere. If the ISCO is located within the ergosphere, it itself serves as the upper boundary of the effective region.
@@
Considering that an ejected plasmoid might be significantly accelerated at the $X$-point and escape to the distance, its emission profile would be markedly different from that of the background plasma fluid. Thus, from an astronomical viewpoint, it is meaningful to investigate the optical appearance of such plasmoid [47, 48]. As the observational precision of the Event Horizon Telescope (EHT) and the GRAVITY collaboration continues to improve [49, 50], it is anticipated that the characteristics of these plasmoids can be detected in the radio and near-infrared bands in the future [51]. This would enable us to identify their energies and origins, and potentially to test energy extraction mechanisms.

### v2

@@ Abstract
Magnetic reconnection within a highly magnetized plasma has been seen as a viable mechanism to extract the energy from a rotating black hole, as it can generate negative energy plasmoids in the ergoregion. For a typical accreting black hole, the ergoregion is filled with bulk plasma plunging from the innermost-stable-circular orbit (ISCO). In this study, we present an analytical study of the energy extraction via magnetic reconnection process in the plunging region. In contrast to the toroidal plasma, where the magnetic field cannot be derived from the MHD scheme, the magnetic field in the plunging plasma was determined by the ideal-MHD condition. We derive the global magnetic field structure in a fast reconnection model, and we read the expressions for the energies of plasmoids ejected from the reconnection region, for general stationary and axisymmetric spacetimes. Then, we demonstrate the behaviors of ejection energies varying with the reconnection locations in the Kerr spacetime, and identify the region where a negative-energy plasmoid can be produced. We find that for a certain magnetization there exists a critical value of the black hole spin, beyond which the energy extraction can occur, and the energy extraction is most efficient for the near-extreme black hole. Moreover, we study the conditions necessary for a plasmoid with positive energy to escape to the infinity, a crucial requirement for effective energy extractions. Considering the escaping conditions, we provide the parameter space in the radius-spin plane in which the energy extraction mechanism is effective.
1 Department of Physics, School of Physical Science and Technology, Ningbo University, Ningbo, Zhejiang 315211, China
2School of Physics, Peking University, No.5 Yiheyuan Rd, Beijing 100871, P.R. China
3Center for High Energy Physics, Peking University, No.5 Yiheyuan Rd, Beijing 100871, P. R. China
@@ 1 Introduction
Beyond the above mechanisms, it has been proposed theoretically that the energy extraction is viable via magnetic reconnections. Magnetized accretion flows around black holes can liberate electromagnetic energy in the regions where magnetic field lines are highly curved and reconnect to discharge their energy [8, 9, 10, 11]. This energy release propels the plasma ejected from the reconnection region. If the released magnetic energy is sufficiently strong, the accelerated plasma can obtain sufficiently negative angular momentum and negative energy, as viewed from the infinity. Koide and Arai proposed the possibility of energy extraction from black holes through magnetic reconnection [12], where they treated the ejected plasma as relativistic adiabatic incompressible ball, without delving into the detailed dynamics of the plasmoids. More recent studies has shown that the plasmoid instability within the reconnection layer and the generation of expelled plasmoids can lead to fast reconnection and fast Alfvén outflows from these layers [13, 14, 15, 16]. Subsequently, Comisso and Asenjo [17] investigated the energy extraction through the plasmoids in a fast reconnection model [18]. They calculated the energy extraction efficiency and power, showing that such an energy extraction rate could be comparable to that of the BZ process, thereby implying its astrophysical significance. Subsequent studies have explored the energy extraction via magnetic reconnections in different types of spacetimes [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
@@ 2.1 The bulk plasma
Strictly speaking, the horizon-scale plasma is nearly collisionless, with the electron free path possibly comparable to the gravitational radius. The intrinsic properties, such as the energy distributions of thermal electrons in both the ejected plasmoids and the bulk plasma, should be described within the framework of Vlasov gas dynamics. However, for simplicity, our current study employs the perfect fluid approximation. The stress-energy tensor of the bulk plasma is given by
@@
where $e=-\alpha T^{t}_{\,\,\,t}$ is called the energy-at-infinity density, and $S^{i}=-\alpha T^{i}_{\,\,\,t}$ is the flux density. In our work, we are interested in a stationary, axisymmetric and geometrically thin fluid configuration. Thus, all the physical quantities are defined over the equatorial plane, neglecting their $\theta$-dependence. The flow is considered to be dominated by the gravity and flow along timelike geodesics, such that $U_{\mu}dx^{\mu}=-Edt+U_{r}(r)dr+Ld\phi$, with $E,L$ being the conserved quantities along the streamline.
@@ 2.3 The global magnetic field structure
Furthermore, it is important to note that the derivation of the global magnetic field mentioned earlier does not explicitly rely on the fluid assumption. What we established is a stationary, axisymmetric, magnetized plasma system characterized by a bulk velocity $U^{\mu}$ at each position. The plasma exhibits sufficiently high conductivity, such that in its rest frame the electrons promptly react and reorganize to neutralize any electric field $E^{\mu}=F^{\mu\nu}U_{\nu}$ that might exist. The timescale associated with this conductivity is shorter than the plunging timescale, ensuring a consistent $E^{\mu}=0$. These conditions allow for the derivation of $B^{\mu}(r,\theta)$. Further details can be found in [38, 47].
@@ 2.4 Energies of ejected plasmoids
During the reconnection process, most of the magnetic energy is converted to the plasma kinetic energy in the current sheet. Thus, within an plasmoid, we can neglect the magnetic energy and only consider the fluid part. On one hand, we take the incompressible sphere approximation, under which the proper thermal quantities of plasma keep constant during the reconnection process [12]. Thus, the proper enthalpy density in the plasmoid is equal to that of the bulk plasma at the $X$-point, and the enthalpy reads $\Delta H=\omega\Delta V$, assuming that the plasmoid has a small proper volume $\Delta V$. On the other hand, the fluid energy-at-infinity density is defined in the normal tetrad, taken to be $e=-\alpha T^{t}_{\,\,\,t}=-\alpha(\omega u^{t}u_{t}+p)$, as has been introduced in Eq. (2.5). The volume of the plasmoid measured by the normal tetrad is $\Delta\hat{V}=\Delta V/\hat{u}^{(t)}=\Delta V/(\alpha u^{t})$, and the energy is $\Delta E=e\Delta\hat{V}$. Based on these, we can define the energy-at-infinity per enthalpy for the ejected plasmoids at the $X$-point as
@@
Note that Eq. (2.4) is analogous to Equation (28) in [17]. The above discussion is quite general and applicable to any stationary and axisymmetric spacetime. According to Eq. (2.4), it is clear that the influence of the spacetime is directly encoded in $\alpha$ and $\beta^{\phi}$, while indirectly encoded in the fluid velocity and the magnetic field angle. The value of $v_{out}$ depends on the reconnection model, which has no association with the background spacetime.
@@ 3 Energy extraction from a Kerr geometry
In what follows we only consider the prograde case. [^2] For the retrograde case, we confirm that the plasmoids have $\epsilon_{\pm}>0$, even under strong magnetizations. Thus, energy extractions via magnetic reconnection cannot occur within the retrograde plasma. In the Kerr geometry, the ISCO is the triple root of the effective potential for radial motion, leading to a simple form of the radial velocity of the plunging geodesics [48],
@@ 3.1 The magnetic field angle
Figure 2: Behaviors of the magnetic field angle in the bulk plasma plunging from the ISCO (Solid lines) in Kerr geometry. We also plot the results for the infalling plasma with $E=1,L=0$ as a comparision (Dashed lines) . Left: $\xi_{B}$ as a function of $r$ for different $a$ (0.8, 0.9, 0.998). Right: The value of $\xi_{B}$ at the event horizon $r_{H}=1+\sqrt{1-a^{2}}$, as a function of $a$.
@@ 3.2 The energy extraction efficiency
Figure 4: The local energy extraction efficiency within the plunging plasma. Top: $\eta$ as a function of the reconnection location, for different values of $a$. The magnetization is fixed at $\sigma_{0}=30$. Bottom left: the maximal value of $\eta$ as a function of the black hole spin. Bottom right: The maximal value of $\eta$ as a function of the magnetization.
In the top panel of Fig. 4, we depict the behavior of the local energy extraction efficiency as a function of the reconnection location within the ISCO. Our findings indicate that for a strong magnetization scenario where the ISCO is situated inside the ergosphere, i.e., $r_{ms}\leq 2$, the energy extraction is inevitably localized near the ISCO. This observation becomes evident when we consider the limit $\sigma_{0}\gg 1$, which yields $v_{out}\approx 1$, $\gamma_{out}\approx\sqrt{(1-{\sf g}^{2})^{3}\sigma_{0}}/(1+{\sf g}^{2})$ with $\sf g$ $=0.49$ for the high magnetization limit. In the vicinity of the ISCO, we have $U^{r},\hat{v}^{(r)}_{s}\rightarrow 0$. Hence, from Eq. (2.4), Eq. (3.7) we can derive
@@
The bottom left panel of Fig. 4 displays the maximum value of $\eta$ under various magnetizations, which depends on the spin of the black hole. For moderate magnetization, such as $\sigma_{0}=3$, only black holes with a rather high spin permit energy extraction. In the case of extreme magnetization, exemplified by $\sigma_{0}=100$, the local energy extraction efficiency exceeds 1 across the range of $a\gtrsim 0.9$. In all these scenarios, the rate of increase of $\eta_{max}$ with respect to $a$ escalates considerably as $a$ approaches its extreme value. To estimate the maximum value of $\eta_{max}$ for a near-extreme black hole, we focus on the near-ISCO region, where the fluid velocity behaves as
@@
Moreover, in the plunging region, the plasma magnetization can be extremely high, surpassing $\sigma_{0}=100$ by several orders of magnitude. For this reason, we present the maximum value of $\eta$ as it scales with magnetization in the bottom right panel of Fig. 4. We observe that once $\sigma_{0}$ exceeds approximately 100, the efficiency becomes nearly a constant. Therefore, at least in the study of energy extraction, we can safely use a typical value of $\sigma_{0}=100$ to represent the cases of extremely high magnetization.
Additionally, we also discuss the case involving the infalling bulk plasma in A.1, where we confirm that energy extraction cannot occur within the infalling bulk plasma.
@@ 4 Summary and Discussions
In this study, we conducted an in-depth study of energy extraction from a Kerr black hole via magnetic reconnections occurring in the plunging region. We initiated our study by considering a general stationary and axisymmetric spacetime, assuming that the plasma fluid in the equatorial plane adheres to the Killing symmetry of the spacetime. The accompanied magnetic field was set to reverse its direction near the equatorial plane, a region conducive to potent magnetic reconnection. Utilizing the ideal-MHD condition, we obtained the global structure of the magnetic field, which fixes the ejection directions of the plasmoids. Our following analysis incorporated a fast reconnection model characterized by strong magnetic pressure, wherein electromagnetic energy is swiftly liberated, resulting in the formation of two high-velocity plasmoids moving in opposite directions out of the current sheet. Subsequently, we derived the analytical form of the energy-at-infinity per enthalpy for the plasmoids.
@@
Furthermore, we investigated the conditions necessary for a plasmoid with positive energy to escape to the infinity, a crucial requirement for the completion of the energy extraction. By scrutinizing the equation of motion for the plasmoid, we pinpointed a critical radius of the $X$-point, beyond which the plasmoid with positive energy can successfully escape. By combining the escaping conditions with the requirement for producing a negative energy plasmoid, we determined the region where energy extraction is effective, as shown in the green region in Fig. 6. Broadly speaking, the lower boundary of the effective region is dictated by the critical radius for the escaping, and the upper boundary is determined by the upper limit for negative energy when the ISCO is situated outside the ergosphere. If the ISCO is located within the ergosphere, it itself serves as the upper boundary of the effective region.
@@
Considering that a plasmoid might be significantly accelerated at the $X$-point and escape to the distance, its emission profile would be markedly different from that of the background plasma fluid. Thus, from an astronomical viewpoint, it is meaningful to investigate the optical appearance of such plasmoid [51, 52]. As the observational precision of the Event Horizon Telescope (EHT) and the GRAVITY collaboration continues to improve [53, 54], it is anticipated that the characteristics of these plasmoids can be detected in the radio and near-infrared bands in the future [55]. This would enable us to identify their energies and origins, and potentially to test energy extraction mechanisms.